Sam owns a business that has $45,000 per month in net cash flows. Assuming
this business can grow by 3% annually in perpetuity and the appropriate discount
rate is 8% what is the current value of this business? (answer is 11,124,000, I just need to know how to do it please and thank you)
Answer #1

Discount rate=   8%  
Growth rate=   3%  
Current Cash flow per month =   45000  
Per year cash flows = 45000*12 months=   540000  
      
Expected cash flow of next year = Current cash flows*(1+growth rate)      
540000*(1+3%)=   556200  
      
Present value of Business as per Constant Growth model= Expected cash flows next year /(Discount Rate - growth rate)      
556200/(8%-3%)=   11124000  
      
So Present Value of business is   $11,124,000.00  
      

(Please thumbs up)